### v1

@@ Abstract
In multi-label emotion classification, particularly for low-resource languages like Arabic, the challenges of class imbalance and label correlation hinder model performance, especially in accurately predicting minority emotions. To address these issues, this study proposes a novel approach that combines stacked embeddings, meta-learning, and a hybrid loss function to enhance multi-label emotion classification for the Arabic language. The study extracts contextual embeddings from three fine-tuned language models—ArabicBERT, MarBERT, and AraBERT—which are then stacked to form enriched embeddings. A meta-learner is trained on these stacked embeddings, and the resulting concatenated representations are provided as input to a Bi-LSTM model, followed by a fully connected neural network for multi-label classification. To further improve performance, a hybrid loss function is introduced, incorporating class weighting, label correlation matrix, and contrastive learning, effectively addressing class imbalances and improving the handling of label correlations. Extensive experiments validate the proposed model’s performance across key metrics such as Precision, Recall, F1-Score, Jaccard Accuracy, and Hamming Loss. The class-wise performance analysis demonstrates the hybrid loss function’s ability to significantly reduce disparities between majority and minority classes, resulting in a more balanced emotion classification. An ablation study highlights the contribution of each component, showing the superiority of the model compared to baseline approaches and other loss functions. This study not only advances multi-label emotion classification for Arabic but also presents a generalizable framework that can be adapted to other languages and domains, providing a significant step forward in addressing the challenges of low-resource emotion classification tasks.
@@ 2.2 Contextual Embeddings with Pre-Trained Language Models
The effectiveness of supervised technologies in EC depends significantly on the features used in classification tasks [29]. Traditional WE techniques include frequency-based methods like Bag of Words (BoW), N-grams, and Term Frequency–Inverse Document Frequency (TF-IDF), as well as neural network-based methods like word2vec [31], Global Vectors for Word Representation (GloVe) [32], and FastText [33]. These earlier models, however, struggled to capture complex semantic dependencies and contextual information, often representing words with single, context-independent vectors [34, 35].
Recent advancements in NLP have led to the development of context-aware Pretrained Language Models (PLMs) such as Bidirectional Encoder Representations from Transformers (BERT) [36]. These models, utilizing transfer learning, have significantly improved performance across various NLP tasks. Unlike earlier WE models, PLMs like BERT excel in capturing complex semantic dependencies and contextual information by using transformer architecture for sub-word learning [36].
@@ 3.1 Dataset Description
Utilizing existing datasets for text emotion classification is common among researchers. However, many online datasets suffer from significant class imbalances, impeding the performance of machine learning (ML) and deep learning (DL) models [64]. Such imbalances often lead models to favor majority classes, resulting in suboptimal predictions for minority classes.
@@
The dataset comprises eleven emotion classes, including anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise, and trust [21]. The dataset was collected using the Twitter API, with search queries aligned to tweets containing emotional words. Annotation reliability was determined using Best Worst Scaling (BWS), with each tweet labeled with ’1’ indicating the presence of the corresponding emotion class and ’0’ indicating its absence. Table 2 provides some sample instances containing Tweets and the multi-label emotion class identified for each Tweet.
@@
To mitigate this challenge, this study proposes the utilization of an optimized loss function specifically designed to address class imbalance issue. By focusing on this aspect, the research aims to enhance the performance of MLEC models on the SemEval-2018 Task 1-Ec-Ar dataset, thereby facilitating more balanced and accurate emotion classification. This approach contributes to the development of more robust ML and DL models and enhances the reliability and applicability of emotion classification in Arabic text.
@@ 3.2 Pre-Processing
The preprocessing of text data is a crucial step in developing robust machine learning and deep learning models for NLP tasks such as MLEC. The preprocessing involved several steps to handle various text formatting issues, particularly focusing on the treatment of punctuation, emojis, emoticons, and normalization of Arabic text. Figure 3 provides a flowchart depicting the steps and their sequence to perform preprocessing of the text.
@@ 3.5 Loss Function
The performance of a deep learning model, especially in the context of multi-label classification tasks, is significantly influenced by the choice of the loss function. For the task of Arabic MLEC, various loss functions have been explored to handle class imbalance and leverage label correlations effectively. This section details the three loss function variations considered: class weights calculation, label correlation matrix, and contrastive learning.
@@ 3.5.1 Class Weights Calculation
Class imbalance is a prevalent issue in multi-label classification, where some classes may have significantly more instances than others as depicted in section 3.1. To address this, class weights are introduced into the loss function to give more importance to minority classes.
@@
  • $y_{c}$ is the ground truth label for class $c$ ($1$ if the instance belongs to class $c$, 00 otherwise).
@@ 3.5.3 Contrastive Learning
Contrastive learning is a technique used to learn representations by comparing similar and dissimilar pairs of instances. In the context of multi-label classification, this approach can be adapted to improve the model’s ability to distinguish between different labels. The contrastive loss function is defined as eq. (5).
@@
These three loss function variations—class weights calculation, label correlation matrix, and contrastive learning—provide robust mechanisms to address the challenges of multi-label emotion classification in imbalanced datasets. Class weights calculation ensures that minority classes receive higher importance during training, the label correlation matrix captures dependencies between labels to improve the prediction of rare classes, and contrastive learning enhances the model’s ability to differentiate between majority and minority classes, leading to improved overall performance.
@@ 3.5.4 Proposed Hybrid loss
    Contrastive Learning (CL): Contrastive Learning helps the model focus on hard-to-classify instances by maximizing the distance between instances of different labels and minimizing the distance between similar ones. This mechanism enhances the model’s ability to differentiate between classes, improving classification accuracy, especially for minority classes.
@@
    Class Weights (CW): Class Weights mitigate the problem of class imbalance by assigning higher weights to minority classes and lower weights to majority classes. This ensures that the model pays more attention to underrepresented classes during training, leading to more balanced classification performance across all labels.
@@ 3.6 Performance Evaluation Measures
To assess the performance of our MLEC model, we used several evaluation metrics. These metrics provide a comprehensive view of the model’s accuracy, ability to identify relevant instances, and its overall effectiveness in handling multi-label classification tasks.
@@ 3.6.1 Precision
Micro-averaged Precision measures the overall accuracy of the positive predictions across all labels. It is defined as the ratio of the sum of true positive predictions to the total number of positive predictions. The formula for micro-averaged precision is given by eq. (6):
@@ 3.6.2 Recall
Micro-averaged Recall measures the model’s overall ability to identify all relevant instances across all labels. It is defined as the ratio of the sum of true positive predictions to the total number of actual positives. The formula for micro-averaged recall is given by eq. (7)
@@ 3.6.3 F1-Score
Micro-averaged F1-Score is the harmonic mean of micro-averaged precision and micro-averaged recall, providing a balance between the two. It is particularly useful for evaluating models on imbalanced datasets. The formula for the micro-averaged F1-score is given by eq. (8):
@@ 3.6.4 Jaccard Accuracy
Jaccard Accuracy measures the average similarity between the predicted and ground truth label sets. It is defined as the size of the intersection divided by the size of the union of the predicted and true label sets. The formula for Jaccard accuracy is given by eq. (9):
@@
where $N$ is the number of instances, $Y^{i}$ is the predicted label set for instance ii, and $Yi$ is the true label set for instance $i$. The numerator $Y^{i}\cap Yi$ represents the size of the intersection of the predicted and true label sets, and the denominator $Y^{i}\cup Yi$ represents the size of their union.
@@ 3.6.5 Hamming Loss
Hamming loss is the fraction of incorrect labels to the total number of labels, providing a measure of how many times an incorrect label is assigned. The formula for Hamming loss is given by eq. (10):
@@
where $N$ is the number of instances, $C$ is the number of classes, $\hat{Y}_{ij}$ is the predicted value for class $j$ of instance $i$, $Y_{ij}$ is the true value for class $j$ of instance $i$, and $I$ is the indicator function that returns $1$ if the predicted and true values are different and 00 otherwise.
These evaluation metrics—precision, recall, F1-score, Jaccard accuracy, and Hamming loss—provide a comprehensive assessment of the model’s performance in multi-label emotion classification for Arabic text, highlighting its strengths and areas for improvement.
@@ 4 Results & Discussion
This section presents the experimental results obtained from evaluating the proposed MLEC model for Arabic text. Extensive experiments are conducted to assess the performance of proposed model across several datasets, applying the hybrid loss function to handle class imbalance and label correlations effectively. The results demonstrate the effectiveness of proposed approach, which combines stacked embeddings extraction, hybrid loss and the meta-learner. The results showcase improvements in key evaluation metrics such as precision, recall, F1-score, Jaccard accuracy, and Hamming loss. Through a detailed analysis of these metrics, we provide insights into the strengths and limitations of our model, highlighting its capability to advance the state-of-the-art in Arabic MLEC.
@@ 4.1 Experimental Results
Table 3 presents the experimental results evaluated against five metrics: Precision, Recall, F1-Score, Jaccard Accuracy, and Hamming Loss. The results compare the baseline method, which uses standard categorical cross-entropy as the loss function, with four variants: Class Weighting (CW), Label Correlation Matrix (LCM), Contrastive Class Weights (CL), and the proposed Hybrid Loss.
@@ 4.1.1 Comparison with State-of-the-Art
In this section, we compare the results of our proposed model with existing state-of-the-art models in Arabic MLEC. Table 4 summarizes the performance of these models using precision, recall, micro & macro F1-scores and accuracy. The proposed model demonstrates superior performance compared to the existing state-of-the-art models across multiple metrics. Notably, the proposed model achieves a Micro-Averaged F1-Score of 0.81 and a Jaccard Accuracy of 0.67, which are significantly higher than those of previously reported models.
@@ 4.1.2 Class-wise Performance
In this section, we provide a detailed analysis of the performance for each class individually, taking into account the number of samples in each class as shown in Table 5. The dataset exhibits class imbalance, a common challenge in multi-label classification tasks. The issue of class imbalance is significantly reduced by the proposed MLEC due to utilization of hybrid loss, which is designed to mitigate its effects but it still has an impact on class-wise performance as shown in The Table 5.
@@ 4.3 Discussion
The experimental results of Table 3 provide some interesting highlights about the performance of the proposed model which are highlighted below:
@@ 4.3.1 Comparative Performance
Comparison of the proposed approach with state-of-the-art indicate that the proposed approach is superior to existing approaches. The improvements in performance can be attributed to several key factors:
@@
  • Contrastive Learning: The use of contrastive learning further refines the model’s ability to distinguish between different classes by learning from both similar and dissimilar instances. This contributes to higher precision and recall values.
@@
As a result, the proposed model significantly outperforms existing models mentioned in Table 4, particularly in terms of F1-Score and Jaccard Accuracy. These improvements underscore the effectiveness of our approach in addressing class imbalance and leveraging label correlations, leading to enhanced classification performance in the domain of Arabic emotion classification.
@@ 5 Conclusion
This study presents a comprehensive approach to improving multi-label emotion classification (MLEC) for the Arabic language by addressing critical challenges such as class imbalance and label correlations. The proposed model integrates stacked embeddings, meta-learning and formulation of a hybrid loss function by integrating class weighting, label correlation matrix, and contrastive learning to effectively tackle these issues. Through rigorous experimentation and performance evaluation, it is evident that the proposed approach significantly enhances the classification performance across various metrics, including Precision, Recall, F1-Score, Jaccard Accuracy, and Hamming Loss. The class-wise performance analysis further demonstrates that the use of hybrid loss successfully reduces the disparity between majority and minority classes, ensuring a balanced and robust model capable of accurately predicting a wide range of emotions. Moreover, the ablation study underscores the contribution of each component of the proposed approach, validating its superiority over baseline and other variant loss functions. The detailed comparison with backbone architectures such as ArabicBERT, MarBERT, and AraBERT confirms the effectiveness of the proposed approach in leveraging these models to achieve state-of-the-art performance in emotion classification tasks. In conclusion, the proposed approach represents a significant advancement in the field of multi-label emotion classification for the Arabic language. It not only addresses the prevalent issues of class imbalance and label correlations but also provides a generalizable framework that can be extended to other languages and domains.
@@
  • Funding
  • Conflict of interest/Competing interests (check journal-specific guidelines for which heading to use)
  • Data availability
  • Code availability
  • Author contribution

### v3

@@ Abstract
In multi-label emotion classification, particularly for low-resource languages like Arabic, the challenges of class imbalance and label correlation hinder model performance, especially in accurately predicting minority emotions. To address these issues, this study proposes a novel approach that combines stacked embeddings, meta-learning, and a hybrid loss function to enhance multi-label emotion classification for the Arabic language. The study extracts contextual embeddings from three fine-tuned language models—ArabicBERT, MarBERT, and AraBERT—which are then stacked to form enriched embeddings. A meta-learner is trained on these stacked embeddings, and the resulting concatenated representations are provided as input to a Bi-LSTM model, followed by a fully connected neural network for multi-label classification. To further improve performance, a hybrid loss function is introduced, incorporating class weighting, label correlation matrix, and contrastive learning, effectively addressing class imbalances and improving the handling of label correlations. Extensive experiments validate the proposed model’s performance across key metrics and a staggering Jaccard accuracy of 0.81, F1-score of 0.67 and a hamming loss of 0.15. The class-wise performance analysis demonstrates the hybrid loss function’s ability to significantly reduce disparities between majority and minority classes, resulting in a more balanced emotion classification. An ablation study highlights the contribution of each component, showing the superiority of the model compared to baseline approaches and other loss functions. This study not only advances multi-label emotion classification for Arabic but also presents a generalizable framework that can be adapted to other languages and domains, providing a significant step forward in addressing the challenges of low-resource emotion classification tasks.
@@ 2.2 Contextual Embeddings with Pre-Trained Language Models
The features employed in classification tasks have a significant impact on how well supervised methods perform in EC [29]. Frequency-based methods like N-grams, Bag of Words, and Term Frequency-Inverse Document Frequency are examples of traditional WE techniques whereas GloVe [31], word2vec [32] and FastText are the neural networks based WE techniques [33]. These earlier models, however, struggled to capture complex semantic dependencies and contextual information, often representing words with single, context-independent vectors [34, 35].
Context-aware Pretrained Language Models (PLMs), like Bidirectional Encoder Representations from Transformers (BERT) [36], have been developed as a result of recent developments in NLP. These models, utilizing transfer learning, have significantly improved performance across various NLP tasks. Unlike earlier WE models, PLMs like BERT excel in capturing complex semantic dependencies and contextual information by using transformer architecture for sub-word learning [36].
@@ 3.1 Dataset Description
Utilizing existing datasets for text based emotion classification is common among researchers. However, many online datasets suffer from significant class imbalances, impeding the performance of supervised learning models [64]. Such imbalances often lead models to favor majority classes, resulting in suboptimal predictions for minority classes.
@@
The dataset comprises eleven emotion classes, including anger, fear, anticipation, joy, disgust,love, pessimism, optimism, trust, surprise and sadness [21]. The dataset was collected using the Twitter API, with search queries aligned to tweets containing emotional words. Annotation reliability was determined using Best Worst Scaling (BWS), with each tweet labeled with ’1’ indicating the presence of the corresponding emotion class and ’0’ indicating its absence. Table 2 provides some sample instances containing Tweets and the multi-label emotion class identified for each Tweet.
@@
This study proposes use of an optimized loss function created especially to address the problem of class imbalance in order to mitigate this challenge. In order to improve MLEC model performance on the SemEval-2018 Task 1-Ec-Ar dataset and enable more accurate and balanced emotion classification, the research will concentrate on this aspect. This approach improves the accuracy and aplicability of emotion classification in Arabic text while also helping to build stronger ML and DL models.
@@ 3.2 Pre-Processing
Text data preprocessing is a critical step in creating robust machine learning models for NLP tasks like MLEC. The preprocessing involved several steps to handle various text formatting issues, particularly focusing on the treatment of punctuation, emojis, emoticons, and normalization of Arabic text. Figure 3 provides a flowchart depicting the steps and their sequence to perform preprocessing of the text.
@@ 3.5 Loss Function
The choice of a loss function has a major impact on a deep learning model’s performance. Different loss functions have been investigated in this study for the Arabic MLEC task in order to effectively handle class imbalance and take advantage of label correlations. This section describes the three loss function variations that were taken into consideration: (i) contrastive learning, (ii) label correlation matrix, and (iii) class weights calculation.
@@ 3.5.1 Class Weights Calculation
In multi-label classification, class imbalance is a common problem where certain classes may have substantially more instances than others, as section 3.1 demonstrates. This is addressed by adding class weights to the loss function, which give more emphasis to minority classes.
@@
  • $y_{c}$ is the ground truth label for class $c$ ($1$ for class $c$, 00 otherwise).
@@ 3.5.3 Contrastive Learning
The process of learning representations by contrasting similar and dissimilar pairs of instances is known as contrastive learning. This method can be modified for multi-label classification to enhance the model’s capacity to discriminate between various labels. The contrastive loss function is defined as eq. (5).
@@
These three loss function variations-class weights calculation, label correlation matrix, and contrastive learning-provide robust mechanisms to address the challenges of multi-label emotion classification in imbalanced datasets. Class weights calculation ensures that minority classes receive higher importance during training, the label correlation matrix captures dependencies between labels to improve the prediction of rare classes, and contrastive learning enhances the model’s ability to differentiate between majority and minority classes, leading to improved overall performance.
@@ 3.5.4 Proposed Hybrid loss
    Contrastive Learning (CL): Contrastive learning maximizes the distance between instances of different labels and minimizes the distance between similar ones, which helps the model concentrate on instances that are difficult to classify. This mechanism improves classification accuracy, particularly for minority classes, by strengthening the model’s capacity to distinguish between classes.
@@
    Class Weights (CW): Class weights help address the issue of class imbalance by giving minority classes more weight and majority classes less. This guarantees that during training the model will focus more on underrepresented classes, resulting in more balanced classification performance across all labels.
@@ 3.6 Performance Evaluation Measures
The efficacy of the MLEC model was assessed using a variety of evaluation metrics. These metrics give an in-depth assessment of the model’s accuracy, ability to find pertinent examples, and overall efficiency especially when it comes to multi-label classification tasks.
@@ 3.6.1 Precision
The overall accuracy of the positive predictions across all labels is measured by the micro-averaged precision. It is calculated by dividing the total number of positive predictions by the sum of true positive predictions as given in eq. (6):
@@ 3.6.2 Recall
The model’s overall capacity to recognize all pertinent instances across all classes is gauged by the micro-averaged recall. It can be defined as the ratio of the total number of real positives to the sum of true positive predictions and is given by eq. (7)
@@ 3.6.3 F1-Score
The harmonic mean of micro-averaged precision and micro-averaged recall, which provides a balance between the two, is the micro-averaged F1-Score. It is especially helpful when assessing models on datasets that are unbalanced and is given by eq. (8):
@@ 3.6.4 Jaccard Accuracy
The average similarity between the ground truth and predicted label sets is measured by Jaccard Accuracy. It is calculated by dividing the intersection’s size by the union of the true and predicted label sets and is given by eq. (9):
@@
where $Yi$ is the true label set for instance $i$, $N$ is the number of instances, and $Y^{i}$ is the predicted label set for instance $ii$. The size of the intersection between the true and predicted label sets is represented by the numerator $Y^{i}\cap Yi$, while the size of their union is represented by the denominator $Y^{i}\cup Yi$.
@@ 3.6.5 Hamming Loss
Hamming loss measures the frequency of incorrect label assignment by dividing the number of incorrect labels by the total number of labels and is given by eq.(10):
@@
where $C$ is the number of classes, $N$ is the number of instances, $\hat{Y}_{ij}$ is the predicted value for class $j$ of instance $i$, and $Y_{ij}$ is the true value for class $j$ of instance $i$. Similarly, $I$ is the indicator function that returns $1$ if the predicted and true values are different and 00 otherwise.
The model’s performance in multi-label emotion classification for Arabic text is thoroughly evaluated using these evaluation metrics, which highlight both the model’s advantages and disadvantages. These metrics include precision, recall, F1-score, Jaccard accuracy, and Hamming loss.
@@ 4 Results & Discussion
The experimental findings from evaluating the proposed MLEC model for Arabic text are presented in this section. Extensive experiments are carried out to evaluate the proposed model’s performance across multiple datasets. The results indicate that the proposed approach, which combines the meta-learner, hybrid loss, and stacked embeddings extraction, is effective. Key evaluation metrics like precision, recall, F1-score, hamming loss and Jaccard accuracy have shown an improvement. Through a thorough examination of these metrics, we shed light on our model’s advantages and disadvantages while emphasizing its potential to push the boundaries of Arabic MLEC.
@@ 4.1 Experimental Results
The experimental results are compared using five metrics: precision, recall, F1-score, jaccard accuracy, and hamming loss (see table 3). Class Weighting (CW), Label Correlation Matrix (LCM), Contrastive Class Weights (CL) as weas well as the proposed Hybrid Loss are the four variations that are compared with the baseline approach, which employs standard categorical cross-entropy as the loss function.
@@ 4.1.1 Comparison with State-of-the-Art
This section presents a comparison between our proposed model’s performance and the state-of-the-art models currently available for Arabic MLEC. The accuracy, precision, recall, micro and macro averaged F1-scores are used to summarize the performance of these models in Table 4. The proposed model outperforms the current state-of-the-art models on a number of different metrics. Notably, the proposed model outperforms previously published models with a Micro-Averaged F1-Score of 0.81 and a Jaccard Accuracy of 0.67.
@@ 4.1.2 Class-wise Performance
The performance for each class is examined in detail, taking into consideration the number of samples in each class as indicated in Table 5. Class imbalance is a prevalent issue in multi-label classification problems, and it is evident in this dataset. The proposed MLEC significantly reduces the problem of class imbalance by using hybrid loss, which aims to mitigate its effects but still influences class-wise performance, as The Table 5 demonstrates.
@@ 4.3 Discussion
Table 3’s experimental results offer a few interesting insights about the proposed model’s performance, which are highlighted below:
@@ 4.3.1 Comparative Performance
It is evident from a comparison with state-of-the-art that the proposed approach is better than existing approaches. The improvements in performance can be attributed to several key factors:
@@
  • Contrastive Learning: By learning from both comparable and dissimilar examples, contrastive learning improves the model’s capacity to discriminate across classes. This leads to higher precision and recall values.
@@
Consequently, the proposed approach performs noticeably better than the current models listed in Table 4, especially when it comes to F1-Score and Jaccard Accuracy. These improvements highlight the effectiveness of our approach for addressing class imbalance while taking benefit of label correlations, which improves classification performance in the Arabic emotion classification domain.
@@ 5 Conclusion
In order to improve multi-label emotion classification (MLEC) for the Arabic language, this work offers a thorough method that addresses important issues such as label correlations and class imbalance. To effectively address these problems, the proposed approach combines stacked embeddings, meta-learning, and the development of a hybrid loss function using a label correlation matrix, contrastive learning, and class weighting. It is clear from extensive testing and performance evaluation that the proposed approach significantly enhances classification performance across a number of metrics, including as precision, recall, F1-score, jaccard accuracy, and hamming loss. The class-wise performance analysis also shows that hybrid loss effectively reduces the gap between majority and minority classes, ensuring a stable and balanced model that can correctly predict a wide range of emotions. Furthermore, the ablation study confirms the advantages of the proposed approach over baseline and alternative variants in terms of loss functions by highlighting the contributions of each component. The effectiveness of the proposed approach for using these models to achieve state-of-the-art performance in emotion classification tasks is confirmed by the thorough comparison with backbone designs like ArabicBERT, MarBERT, and AraBERT. Conclusively, the proposed approach is an important advancement in the area of Arabic multi-label emotion classification. It offers a generalizable framework that may be applied to different languages and topics in addition to addressing the common problems of class imbalance and label correlations.
@@
  • Funding: This research was funded by the Chinese Academy of Sciences President’s International Fellowship Initiative, Grant No. 2024PVB0036.
  • Conflict of Interest: The authors declare no conflicts of interest.
  • Code Availability: The code for the model and experiments is accessible via GitHub at https://github.com/nisarahmedrana/ArabicMLEC.
  • Data Availability: The dataset used in this study is available in the same GitHub repository: https://github.com/nisarahmedrana/ArabicMLEC.
  • Author Contributions: M.A.A. led the research and contributed to all phases, including conceptualization, methodology, analysis, and manuscript preparation. W.J. performed analysis and write up of the manuscript. N.A. was responsible for the design, development, modeling, and manuscript writing. M.I.Z. performed the coding and implementation of the proposed method. L.Y., H.H., W.S., and X.L. provided support in data curation, analysis, and validation. All authors reviewed the manuscript and provided valuable feedback.